# Efficient Inference in First Passage Time Models

### v1

@@ 4.2 Integrated Single-stage Model
The single-stage model uses $X(0)=x_{0}$ deterministically. If, instead, $X(0)$ is random, then the corresponding FPTDs and NPDs can be obtained by taking the expectations of (9), (10), and (11) over the distribution of $X(0)$. We call this the integrated single-stage model. If the distribution of $X(0)$ has a probability density function $q_{0}$ supported on $(a_{2},a_{1})$, then FPTDs and NPDs for the integrated single-stage model are
$f_{u}^{\text{int}}(t;\mu,\sigma,\mathcal{B},q_{0})$ $=\int_{a_{2}}^{a_{1}}f_{u}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})q_{0}(x_{0})\,\mathrm{d}x_{0},$ (12)
$f_{\ell}^{\text{int}}(t;\mu,\sigma,\mathcal{B},q_{0})$ $=\int_{a_{2}}^{a_{1}}f_{\ell}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})q_{0}(x_{0})\,\mathrm{d}x_{0},$
$q^{\text{int}}(x;\mu,\sigma,\mathcal{B},q_{0})$ $=\int_{a_{2}}^{a_{1}}q^{\text{single}}(x;\mu,\sigma,\mathcal{B},x_{0})q_{0}(x_{0})\,\mathrm{d}x_{0}.$
An illustration of the integrated single-stage model is shown in Figure 2(b).
Below, we use the same definitions and formulas in equation (12) when $q_{0}$ is a sub-probability density function. This usage forms the basis of our multi-stage model below. When $X(0)$ is a discrete random variable, we abuse notation and allow $q_{0}=\sum_{j=1}^{J}w_{0j}\delta_{x_{0j}}$ to be a weighted sum of point masses, where $(w_{01},\ldots,w_{0J})$ is a probability mass function on $(x_{01},\ldots,x_{0J})$ and $\delta$ is the Dirac delta generalized function. For instance, when $q_{0}=\delta_{x_{0}}$ we have $f_{u}^{\text{int}}(t;\mu,\sigma,\mathcal{B},q_{0})=f_{u}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})$ and similarly for $f_{\ell}^{\text{int}}$ and $q^{\text{int}}$, so that a single-stage model can be viewed as a special case of the integrated single-stage model.
@@ 4.3 Multi-stage Model
stopped by a continuous piecewise linear upper boundary $u(t)$, a continuous piecewise linear lower boundary $\ell(t)$ and a vertical boundary $t=T_{\text{end}}$, where $\mu(t)$ is a piecewise constant drift function and $\sigma(t)$ is a piecewise constant diffusion function. Let $0<t_{1}<\cdots<t_{d-1}<T_{\text{end}}$ be the $d-1$ segment points for $\mu(t),\sigma(t),u(t)$ and $\ell(t)$ prior to $T_{\text{end}}$, and define $t_{0}=0$ and $t_{d}=T_{\text{end}}$. Then $\mu(t),\sigma(t),u(t)$ and $\ell(t)$ can be defined as
$\mu(t)&=\mu_{k}\\ \sigma(t)&=\sigma_{k}\\ u(t)&=a_{k1}+b_{k1}(t-t_{k-1})\\ \ell(t)&=a_{k2}+b_{k2}(t-t_{k-1})\end{aligned}\right\}\text{when }t_{k-1}<t\leq t_{k}$
for $1\leq k\leq d$. Note that we must have $a_{k1}+b_{k1}(t_{k}-t_{k-1})=a_{k+1,1}$ and $a_{k2}+b_{k2}(t_{k}-t_{k-1})=a_{k+1,2}$ for $11\leq k\leq d-$ 1 to ensure the continuity of $u(t)$ and $\ell(t)$. We denote the process during time interval $\left(t_{k-1},t_{k}\right]$ to be the $k$-th (integrated single) stage, which has a drift coefficient $\mu_{k}$, diffusion coefficient $\sigma_{k}$, and boundary parameters $\mathcal{B}_{k}=\left(a_{k1},b_{k1},a_{k2},b_{k2},T_{k}=t_{k}-t_{k-1}\right)$, for $1\leq k\leq d$. By the Markov property, the NPD $q^{\text{int}}$ in equation (12) from the $(k-1)$-st stage becomes the initial position density $q_{0}$ of the $k$-th stage. Since the process could have hit the upper or lower boundary in an earlier stage, the NPD is a sub-probability density function, which is why we allowed $q_{0}$ to be a sub-probability density function in the definition of the integrated single-stage model.
@@
Figure 3: An illustration of the sequential computation from the $k$-th stage to the $(k+1)$-th stage in a multi-stage model. The $k$-th stage (depicted on the left side of the figure) is an integrated single-stage model, where the density of its starting position is given by $q_{k-1}(x)$ (represented by the grey curve). Within this stage, the functions $f_{u}(t)\mathbbm{1}_{(t_{k-1},t_{k}]}(t)$ (red curve), $f_{\ell}(t)\mathbbm{1}_{(t_{k-1},t_{k}]}(t)$ (purple curve), and $q_{k}(x)$ (green curve) are obtained by integrating the formulas (9), (10), and (11) over $q_{k-1}$, respectively. The resulting NPD $q_{k}(x)$ then serves as the starting position distribution for the $(k+1)$-th stage, facilitating the subsequent computation (depicted on the right side of the figure).
@@
for $k=1$ to $d$ do
$f_{u}(t)&=f_{u}^{\text{int}}(t-t_{k-1};\mu_{k},\sigma_{k},\mathcal{B}_{k},q_{k-1})\\ f_{\ell}(t)&=f_{\ell}^{\text{int}}(t-t_{k-1};\mu_{k},\sigma_{k},\mathcal{B}_{k},q_{k-1})\\ \end{aligned}\right\}\text{ for }t\in(t_{k-1},t_{k}]$
@@
For multi-stage models, Algorithm 1 is exact in the sense that the only error is the numerical error. We note, however, that equations (6) and (7) are singular at $t=0$, which can lead to numerical instability in Algorithm 1 if the observation time $t\in(t_{k-1},t_{k}]$ is very close to $t_{k-1}$. In practical scenarios, this issue can usually be avoided via data post-processing that slightly inflates the value of any $t$ that happens too close to the beginning of a stage. This will have negligible impact on parameter estimation, particularly in cognitive psychology applications where any numerical corrections are going to lie well within the measurement error of participant reaction times.
@@ 4.4 Approximating with the Multi-stage Model
For GDDMs of the same form as the multi-stage model, but with boundaries that are not piecewise linear, it is natural to consider approximating the GDDM with a multi-stage model by approximating the boundaries $u(t)$ and $\ell(t)$ with continuous piecewise linear functions. Choosing time points $0=t_{0}<t_{1}<\dotsb<t_{d}=T_{\text{end}}$ that include any times when $\mu(t)$ or $\sigma(t)$ changes values, we can define $\bar{u}(t)$ and ¯$\bar{\ell}(t)$ to be the continuous piecewise linear functions that are linear on each interval $(t_{k-1},t_{k}]$ and for which $\bar{u}(t_{k})=u(t_{k})$ and ¯$\bar{\ell}(t_{k})=\ell(t_{k})$ for each $k$. The original process with these new boundaries gives a multi-stage model that approximates the original GDDM. The FPTDs and NPDs of the approximating multi-stage model can be computed with Algorithm 1 and used as approximations of the FPTDs and NPDs for the original GDDM.
Define $\bar{\tau}=\inf\{t>0:X(t)\not\in(\bar{\ell}(t),\bar{u}(t))\}$ to be the hitting time of the process to the approximating boundaries. As the approximation gets increasingly fine, meaning $\max_{k}|t_{k}-t_{k-1}|\to 0$, then it is straightforward to show that $\bar{\tau}\to\tau$ with probability one under mild regularity conditions on the boundaries (see the proof in Appendix A). Showing that the FPTDs also converge is beyond the scope of this paper and presumably requires additional regularity assumptions on the boundaries, such as continuous differentiability. We provide empirical evidence in the examples below that this method of approximation works well.
@@ 4.5 Transformations
~$\widetilde{\ell}(s)$ $=\psi(\ell(\gamma^{-1}(s)),\gamma^{-1}(s)),$
@@
$f_{c}(t)=\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{P}(\tau\leq t,C=c)=\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{P}(\widetilde{\tau}\leq\gamma(t),\widetilde{C}=c)=\gamma^{\prime}(t)\widetilde{f}_{c}(\gamma(t))$ (14)
@@
$q(x)$ $=\frac{\,\mathrm{d}}{\,\mathrm{d}x}\mathbb{P}(X(T)\leq x,\tau>T)=\frac{\,\mathrm{d}}{\,\mathrm{d}x}\mathbb{P}(\widetilde{X}(\gamma(T))\leq\psi(x,T),\widetilde{\tau}>\gamma(T))$ (15)
@@
A particularly simple $\widetilde{X}$ is standard Brownian motion, perhaps with a random starting point, i.e., equation (2) with $\mu(x,t)\equiv 0$ and $\sigma(x,t)\equiv 1$. A wide range of commonly used SDEs can be transformed into standard Brownian motion in this way, including geometric Brownian motion, the Ornstein-Uhlenbeck (O-U) process, and the Brownian bridge, among others. The question of whether a diffusion process can be transformed to a Brownian motion through transformation of the type in equation (13) was raised by Kolmogorov in [33]. Cherkasov gave a constructive proof of the necessary and sufficient condition for the existence of such a transformation in [10], which is now commonly referred as the Cherkasov condition in the literature. This was recast into a simpler form in [53], and was later used in [52] to obtain the one-sided first passage time density through the integral equation method. We present the Cherkasov condition in [53] in Appendix B, and instead of focusing on the condition itself, we provide specific examples of the transformation from diffusion processes to Brownian motion that are common in practical applications.
@@
    where $\theta\neq 0,\lambda\in\mathbb{R},\sigma>0$ are constants. When $\theta>0$ it models the “mean-reverting” phenomenon as the process tends towards $\lambda$ over time, and it will converge to its stationary distribution $\mathcal{N}(\lambda,\frac{\sigma^{2}}{2\theta})$ as $t$ goes to infinity. This model has a long history in cognitive science [64], where it is called the leaky competing accumulator model of decision making. This model can flexibly capture evidence accumulation and neural dynamics that are either leaky (for $\theta<0$, such that earlier bits of evidence are disregarded relative to later bits, due to mean reversion), or attractive (for $\theta>0$, where early evidence dominates); see [9].
@@
    $\widetilde{X}(s)=\psi(X(\gamma^{-1}(s)),\gamma^{-1}(s))=\frac{1}{\sigma}X(0)+\widetilde{W}(s)$
@@
    In Section 5.3, we use this transformation and adopt the procedures in this section to compute the FPTDs of the O-U process.
Additionally, for parametric models like the O-U process, we can also study their “multi-stage” versions with piecewise constant parameters, as well as “attentional” versions, where the values of these piecewise constant parameters and the stage durations are influenced by attention mechanisms. Our Algorithm 1 naturally admits generalization to these models, in which case the breakpoints between stages will contain both the saccade times and the segment points for approximating the nonlinear transformed boundaries.
@@ 4.6 Algorithm Summary
4.7 Fast Implementation
In this section, we discuss some practical aspects of our algorithm that contribute to its fast execution, including important details about numerical integration and function representation that are glossed over in the description of Algorithm 1.
  • If we have multiple observations for the same GDDM, then one pass of Algorithm 1 can be used to evaluate the likelihood for all observations for this GDDM, thereby eliminating redundant computations.
  • To compute the likelihood given a dataset, we simply iterate Algorithm 1 over all data. The computation of each trial in a dataset is independent of each other, rendering the problem “embarrassingly parallelizable” across trials — we can distribute the dataset over multiple threads, and since no communication is needed between the sub-tasks, the computational time should ideally be inversely proportional to the number of threads.
  • All numerical integrations in equation (12) as used by Algorithm 1 are performed with Gauss-Legendre quadrature (see Chapter 5.5 of [56] for an introduction). This approach offers higher accuracy with fewer integrand evaluations compared to alternative methods, such as the trapezoidal rule. We specifically benefit from this in our method because our integrands consist of (truncated) infinite series, which are expensive to evaluate. It also means that each $q_{k}$ within the algorithm needs only to be represented at a small number of pre-determined quadrature points. Details of the implementation of Algorithm 1 regarding quadrature are shown in Appendix C.
In our program, we implement the core functionality in Python, with performance-critical parts written using Cython [4], leveraging both the flexibility of Python and the high efficiency of C. OpenMP parallelization is supported to speed up the likelihood computation for large datasets. Our code, along with usage examples—including numerical examples presented in Section 5—will be released as an open-source Python package upon publication for reproducibility. Additionally, our codes will be integrated into the HSSM package [19] for hierarchical Bayesian inference of GDDMs.

### v2

@@ 4.2 Integrated Single-stage Model
The single-stage model uses $X(0)=x_{0}$ deterministically. If, instead, $X(0)$ is random, then the corresponding FPTDs and NPDs can be obtained by taking the expectations of (4.1), (4.1), and (4.1) over the distribution of $X(0)$. We call this the integrated single-stage model. If the distribution of $X(0)$ has a probability density function $q_{0}$ supported on $(a_{2},a_{1})$, then FPTDs and NPDs for the integrated single-stage model are
$f_{u}^{\text{int}}(t;\mu,\sigma,\mathcal{B},q_{0})=$ (12)
$\quad\int_{a_{2}}^{a_{1}}f_{u}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})q_{0}(x_{0})\,\mathrm{d}x_{0},$
$f_{\ell}^{\text{int}}(t;\mu,\sigma,\mathcal{B},q_{0})=$
$\quad\int_{a_{2}}^{a_{1}}f_{\ell}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})q_{0}(x_{0})\,\mathrm{d}x_{0},$
$q^{\text{int}}(x;\mu,\sigma,\mathcal{B},q_{0})=$
$\quad\int_{a_{2}}^{a_{1}}q^{\text{single}}(x;\mu,\sigma,\mathcal{B},x_{0})q_{0}(x_{0})\,\mathrm{d}x_{0}.$
An illustration of the integrated single-stage model when $X(0)$ has a probability density function is shown in Figure 2(b). When $X(0)$ is a discrete random variable, we abuse notation and allow $q_{0}=\sum_{j=1}^{J}w_{0j}\delta_{x_{0j}}$ to be a weighted sum of point masses, where $(w_{01},\ldots,w_{0J})$ is a probability mass function on $(x_{01},\ldots,x_{0J})$ and $\delta$ is the Dirac delta generalized function, so that Eq.(12) still holds formally. For instance, when $q_{0}=\delta_{x_{0}}$ we have $f_{u}^{\text{int}}(t;\mu,\sigma,\mathcal{B},q_{0})=f_{u}^{\text{single}}(t;\mu,\sigma,\mathcal{B},x_{0})$ and similarly for $f_{\ell}^{\text{int}}$ and $q^{\text{int}}$, so that a single-stage model can be viewed as a special case of the integrated single-stage model. We use the same definitions and formulas in equation (12) when $q_{0}$ is a sub-probability density function, and this usage forms the basis of our multi-stage model below.
@@ 4.3 Multi-stage Model
stopped by a continuous piecewise linear upper boundary $u(t)$, a continuous piecewise linear lower boundary $\ell(t)$ and a vertical boundary $t=T_{\text{end}}$, where $\mu(t)$ is a piecewise constant drift function and $\sigma(t)$ is a piecewise constant diffusion function. Let $0<t_{1}<\cdots<t_{d-1}<T_{\text{end}}$ be the d1d-1 segment points for $\mu(t),\sigma(t),u(t)$ and $\ell(t)$ prior to $T_{\text{end}}$, and define $t_{0}=0$ and $t_{d}=T_{\text{end}}$. Then $\mu(t),\sigma(t),u(t)$ and $\ell(t)$ can be defined as
$\left.\begin{aligned} \mu(t)&=\mu_{k}\\ \sigma(t)&=\sigma_{k}\\ u(t)&=a_{k1}+b_{k1}(t-t_{k-1})\\ \ell(t)&=a_{k2}+b_{k2}(t-t_{k-1})\end{aligned}\right\}\text{when }t_{k-1}<t\leq t_{k}$
for $1\leq k\leq d$. Note that we must have $a_{k1}+b_{k1}(t_{k}-t_{k-1})=a_{k+1,1}$ and $a_{k2}+b_{k2}(t_{k}-t_{k-1})=a_{k+1,2}$ for $1\leq k\leq d-1$ to ensure the continuity of $u(t)$ and $\ell(t)$. We denote the process during time interval $\left(t_{k-1},t_{k}\right]$ to be the $k$-th (integrated single) stage, which has a drift coefficient $\mu_{k}$, diffusion coefficient $\sigma_{k}$, and boundary parameters $\mathcal{B}_{k}=\left(a_{k1},b_{k1},a_{k2},b_{k2},T_{k}=t_{k}-t_{k-1}\right)$, for $1\leq k\leq d$. By the Markov property, the NPD $q^{\text{int}}$ in equation (12) from the (k1)(k-1)-st stage becomes the initial position density $q_{0}$ of the $k$-th stage. Since the process could have hit the upper or lower boundary in an earlier stage, the NPD is a sub-probability density function, which is why we allowed $q_{0}$ to be a sub-probability density function in the definition of the integrated single-stage model.
@@
Figure 3: An illustration of the sequential computation from the $k$-th stage to the $(k+1)$-th stage in a multi-stage model. The $k$-th stage (depicted on the left side of the figure) is an integrated single-stage model, where the density of its starting position is given by $q_{k-1}(x)$ (represented by the grey curve). Within this stage, the functions $f_{u}(t)\mathbbm{1}_{(t_{k-1},t_{k}]}(t)$ (red curve), $f_{\ell}(t)\mathbbm{1}_{(t_{k-1},t_{k}]}(t)$ (purple curve), and $q_{k}(x)$ (green curve) are obtained by integrating the formulas (4.1), (4.1), and (4.1) over $q_{k-1}$, respectively. The resulting NPD $q_{k}(x)$ then serves as the starting position distribution for the $(k+1)$-th stage, facilitating the subsequent computation (depicted on the right side of the figure).
@@
1exfor $k=1$ to $d$ do
$\left.\begin{aligned} f_{u}(t)&=f_{u}^{\text{int}}(t-t_{k-1};\mu_{k},\sigma_{k},\mathcal{B}_{k},q_{k-1})\\ f_{\ell}(t)&=f_{\ell}^{\text{int}}(t-t_{k-1};\mu_{k},\sigma_{k},\mathcal{B}_{k},q_{k-1})\\ \end{aligned}\right\}\text{ for }t\in(t_{k-1},t_{k}]$
@@
For multi-stage models, Algorithm 1 is exact in the sense that the only error is the numerical error. We note, however, that equations (6) and (7) are singular at $t=0$, which can lead to numerical instability in Algorithm 1 if the observation time $t\in(t_{k-1},t_{k}]$ is very close to $t_{k-1}$. In practical scenarios, this issue can usually be avoided via data post-processing that slightly inflates the value of any $t$ that happens too close to the beginning of a stage. This will have negligible impact on parameter estimation, particularly in cognitive scientific applications where any numerical corrections are going to lie well within the measurement error of participant reaction times.
@@ 4.4 Approximating with the Multi-stage Model
For GDDMs of the same form as the multi-stage model, but with boundaries that are not piecewise linear, it is natural to consider approximating the GDDM with a multi-stage model by approximating the boundaries $u(t)$ and $\ell(t)$ with continuous piecewise linear functions. Choosing time points $0=t_{0}<t_{1}<\dotsb<t_{d}=T_{\text{end}}$ that include any times when $\mu(t)$ or $\sigma(t)$ changes values, we can define $\overline{u}(t)$ and $\overline{\ell}(t)$ to be the continuous piecewise linear functions that are linear on each interval $(t_{k-1},t_{k}]$ and for which $\overline{u}(t_{k})=u(t_{k})$ and $\overline{\ell}(t_{k})=\ell(t_{k})$ for each $k$. The original process with these new boundaries gives a multi-stage model that approximates the original GDDM. The FPTDs and NPDs of the approximating multi-stage model can be computed with Algorithm 1 and used as approximations of the FPTDs and NPDs for the original GDDM.
Define $\overline{\tau}=\inf\{t>0:X(t)\not\in(\overline{\ell}(t),\overline{u}(t))\}$ to be the hitting time of the process to the approximating boundaries. As the approximation gets increasingly fine, meaning $\max_{k}|t_{k}-t_{k-1}|\to 0$, then it is straightforward to show that $\overline{\tau}\to\tau$ with probability one under mild regularity conditions on the boundaries (see the proof in Appendix A.1). Showing that the FPTDs also converge is beyond the scope of this paper and presumably requires additional regularity assumptions on the boundaries, such as continuous differentiability. We demonstrate in Sections 5.2 and 5.3 that this approximation method can perform remarkably well in practice.
@@ 4.5 Transformations
$\widetilde{\ell}(s)$ $=\psi(\ell(\gamma^{-1}(s)),\gamma^{-1}(s)),$
@@
$f_{c}(t)=$ $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{P}(\tau\leq t,C=c)$ (14)
$=$ $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{P}(\widetilde{\tau}\leq\gamma(t),\widetilde{C}=c)$
$=$ $\gamma^{\prime}(t)\widetilde{f}_{c}(\gamma(t))$
@@
$q(x)$ $=\frac{\,\mathrm{d}}{\,\mathrm{d}x}\mathbb{P}(X(T)\leq x,\tau>T)$ (15)
$=\frac{\,\mathrm{d}}{\,\mathrm{d}x}\mathbb{P}(\widetilde{X}(\gamma(T))\leq\psi(x,T),\widetilde{\tau}>\gamma(T))$
@@
A particularly simple $\widetilde{X}$ is standard Brownian motion, perhaps with a random starting point, i.e., equation (2) with $\mu(x,t)\equiv 0$ and $\sigma(x,t)\equiv 1$. A wide range of commonly used SDEs can be transformed into standard Brownian motion in this way, including geometric Brownian motion, the Ornstein-Uhlenbeck (O-U) process, and the Brownian bridge, among others. The question of whether a diffusion process can be transformed to a Brownian motion through transformation of the type in equation (13) was raised by Kolmogorov in [28]. Cherkasov gave a constructive proof of the necessary and sufficient condition for the existence of such a transformation in [7], which is now commonly referred as the Cherkasov condition in the literature. This was recast into a simpler form in [47], and was later used in [46] to obtain the one-sided first passage time density through the integral equation method. We present the Cherkasov condition in [47] in Appendix A.2, and instead of focusing on the condition itself, we provide specific examples of the transformation from diffusion processes to Brownian motion that are common in practical applications.
@@
where $\theta\neq 0,\lambda\in\mathbb{R},\sigma>0$ are constants. When $\theta>0$ it models the “mean-reverting” phenomenon as the process tends towards $\lambda$ over time, and it will converge to its stationary distribution $\mathcal{N}(\lambda,\frac{\sigma^{2}}{2\theta})$ as $t$ goes to infinity. This process is used, for instance, in physics to model the velocity of a particle under friction, in finance to capture mean-reverting behavior of interest rates and volatility [18], and in cognitive science to model evidence accumulation and neural dynamics [57, 6].
@@
$\widetilde{X}(s)=\psi(X(\gamma^{-1}(s)),\gamma^{-1}(s))=\frac{X(0)}{\sigma}+\widetilde{W}(s)$
@@
In Section 5.3, we use this transformation and adopt the procedures in this section to compute the FPTDs of the O-U process. Additionally, for parametric models like the O-U process, we can also study their “multi-stage” versions with piecewise constant parameters. Our Algorithm 1 naturally admits generalization to these models, where the stage breakpoints comprise both the original parameter breakpoints and the segmentation points introduced to approximate the nonlinear transformed boundaries.
@@ 4.6 Algorithm Summary and Implementation
In our implementation, all numerical integrations in equation (12) as used by Algorithm 1 are performed with Gauss-Legendre quadrature (see Chapter 5.5 of [49] for an introduction). This approach offers higher accuracy with fewer integrand evaluations compared to alternative methods, such as the trapezoidal rule. We specifically benefit from this in our method because our integrands consist of (truncated) infinite series, which are expensive to evaluate. It also means that each $q_{k}$ within the algorithm needs only to be represented at a small number of predetermined quadrature points. The choice of quadrature order provides a natural trade-off between speed and accuracy: higher orders increase the number of quadrature points and hence the computational cost, but also yield greater precision. Details of the implementation of Algorithm 1 regarding quadrature are in Appendix A.3.1.
There are several other important practical aspects of Algorithm 1 that contribute to its efficiency, which we discuss in detail in Appendix A.3. In our implementation, the core functionality is written in Python, while performance-critical components are optimized using Cython [2], leveraging both the flexibility of Python and the high efficiency of C. OpenMP parallelization is supported to compute likelihoods for different trials independently, thereby accelerating the computation on large datasets. Our code, along with usage examples, including the numerical examples presented in Section 5, will be released as an open-source Python package upon publication for reproducibility. Additionally, our code will be integrated into the HSSM package [15] for hierarchical Bayesian inference of GDDMs.